OpenStudy (anonymous):

I keep getting turned around with these types of problems. Factor x^3 -516

jimthompson5910 (jim_thompson5910):

are you sure it's not 512 instead of 516?

OpenStudy (anonymous):

let me check

jimthompson5910 (jim_thompson5910):

or it possibly could be 216

OpenStudy (anonymous):

Sorry i wrote it down wrong, it is 512

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

notice how 8^3 = 512

jimthompson5910 (jim_thompson5910):

so x^3 - 512 is really the same as x^3 - 8^3

jimthompson5910 (jim_thompson5910):

x^3 - 8^3 is a difference of cubes, so use the rule a^3 - b^3 = (a-b)(a^2 + ab + b^2) to factor

OpenStudy (anonymous):

okay I'll go back and try

OpenStudy (anonymous):

is it(x-3)(x^2-3x-64)?

jimthompson5910 (jim_thompson5910):

how are you getting 3?

jimthompson5910 (jim_thompson5910):

in this case a = x b = 8

OpenStudy (anonymous):

from x^3

jimthompson5910 (jim_thompson5910):